OpenStudy (anonymous):

How do you solve this? \[6x ^{2} + 5x \ge 4\]

OpenStudy (anonymous):

start by subtracting 4 from both sides

OpenStudy (anonymous):

then solve for both x

OpenStudy (anonymous):

but the answer says \[\left( - \infty, - \frac{ 4 }{ 3 } \right) \cup \left( \frac{ 1 }{ 2 }, \infty \right)\] what does that mean?

OpenStudy (anonymous):

it means any number between negative infinity and -4/3 And any number between 1/2 and infinity is an answer to your question

OpenStudy (anonymous):

so i just put \[-\frac{ 4 }{ 3 } \] and \[\frac{ 1 }{ 2 }\] will that be ok to put?

OpenStudy (anonymous):

unless you have x <= -4/3 and x >= 1/2 that answer is unacceptable

OpenStudy (anonymous):

ok so i got that so I just write the one with infinity?